OpenStudy (anonymous):

Mrs. Jones Algebra 2 class scored very well on yesterday’s quiz. With one exception, everyone received an A. Within how many standard deviations from the mean do all the quiz grades fall? 91, 92, 94, 88, 96, 99, 91, 93, 94, 97, 95, 97 A. 2 B. 1 C. 3 D. 4

OpenStudy (ageta):

Highest Value = 99 Lowest Value = 88 Number of observations = 12 First calculate the mean and standard deviation, Mean = 93.9 , SD = 3.12 (From analysis) (99-93.9)/3.12 = 1.6 (88 - 93.9)/3.12 = -1.89 all the quiz grades fall within 2 SD from the mean.

OpenStudy (anonymous):

Calculate The Mean First. \[93.9 = 94 \]Now Minus It From Each Value And Square It. So It Should Be . . . \[(91 -94)^{2} = (-3 ^{2})= 9\]

OpenStudy (ageta):

(99-93.9)/3.1 = 5.1/3.1 = 2 (within two standard deviations from mean)

OpenStudy (anonymous):

\[\sqrt{(9+4+0+36+4+25+9+1+0+9+1+9\div12}\]